The hypotenuse of a right triangle is 37 cm. long one leg is 23 cm. longer than the other. Find the length of the legs

Answer by harpazo(655)

I will set up your equation and you can solve for x.
Ready?
The hypotenuse is 37.
One leg is x and the other leg is x + 23.
Use (leg)^2 + (leg)^2 = (hypotenuse)^2

x^2 + (x + 23)^2 = (37)^2
Can you finish?